Grade 11Mechanics

A body is released from the top a tower of height h meter.It takes T seconds to reach the ground.Where is the ball at time T/2 seconds?

Ayush kumar
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear Ayush
 
 

the acceleration of the ball will be g. Initial velocity will be 0. 

 

in T sec. body travels h mts.  

 

 by applying equations of motion we get

 

                                         s= ut +1/2gT2

 

                                         h = 1/2gT2                                          ------[1]

 

in T/2 sec                      h= 1/2gT2/4                                   -------[2]

 

from [1] and [2] we get h=h/4 distance from point of release.

 

therefore distance from ground is h-h/4 =3h/4

 

Regards

Arun(askIITians forum expert)